This server now only holds all the Faulkes Telescope North and Faulkes Telescope South data. Those requiring access to Liverpool Telescope data should see the LT data archive.
- 2013-05-21: The failed Enclosure UPS at FTN has been repaired and tested and the telescope is back in action.
- 2013-05-14: FTS has been returned to action with a recoated primary mirror and new Spectral camera (fs03). Merope (em03) has an electronic problem and will be down for at least a month.
- 2013-05-02: FTN has a failed Enclosure UPS and will be off-sky until the week beginning 20th May.
- 2013-04-02: FTS will be off-sky for maintainance and primary mirror recoating from 13th April until the 25th April inclusive.
- 2013-04-01: The new 2013A semester starts today.